The measure of angle A is four more than one-third the measure of its complement. What is the measure of angle A?
I think that this question is a little funny.

Answer by Alan3354(69443)

The measure of angle A is four more than one-third the measure of its complement. What is the measure of angle A?
-----------------
C = complement
A = C/3 + 4
A + C = 90
C/3 + 4 + C = 90
C + 12 + 3C = 270
4C = 258
C = 64.5 degs
A = 25.5 degs